OpenStudy (anonymous):

QUESTION is the inverse of something the same as putting it in a fraction under one?

OpenStudy (lgbasallote):

yes. for example, the inverse of 2 is 1/2

OpenStudy (anonymous):

ok thanks, and isnt the inverse the same thing is you switch the x and y?

OpenStudy (precal):

Andrew93 that is the procedure for finding the inverse function

OpenStudy (lgbasallote):

that's a different story y= x + 2 the inverse of this is not 1/(x+2) the inverse is actually y = x - 2

OpenStudy (lgbasallote):

but if it's just a number, then the inverse is one over that number

OpenStudy (anonymous):